OpenStudy (anonymous):

y=2cos(2pix) Amplitude___ period______ horizontal shift_______ vertical shift_________

OpenStudy (anonymous):

http://www.purplemath.com/modules/grphtrig.htm

OpenStudy (anonymous):

that's not helping I need the answers

OpenStudy (ybarrap):

$$ y=A\cos(\cfrac{2\pi}{T}x+\phi)+b $$ Where A is the amplitude, T is the period, \(\phi\) is the horizontal shift and b is the vertical shift. I'll help with the period. For your equation, $$ \cfrac{2\pi}{T}=2\pi\\ $$ So T=1. A, \(\phi\) and b you can do easily by inspection. Hope this helps.

OpenStudy (anonymous):

amp is 1 horz and vert shifts are 0?

OpenStudy (ybarrap):

amp is not 1 but you've got the horz and vert shifts right

OpenStudy (anonymous):

@mathmale @Mertsj @ranga can any of you help me?

OpenStudy (mertsj):

That amplitude is the coefficient of the function. In this case it is 2cos so the amplitude is 2

OpenStudy (mertsj):

The coefficient of x is 2pi The coefficient of x = 2pi/period So: 2pi = 2pi/period Solve that for the period

OpenStudy (anonymous):

@Mertsj is the period 1?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

And the shifts arer both 0? @Mertsj

OpenStudy (mertsj):

yes